A tank contains a pool of mercury 0.30m deep,covered with a layer of water that is 1.2m deep. The density of water and Mercury are 1 and 13.6 in g/cm^3. Find the pressure exerted by the two layers at the bottom of the tank.
 
 

Grade:9

1 Answers

Arun
6 years ago
Dear Satvik
 
First find the pressure at the top of the mercury pool.
For a point below the surface of the mercury this may be regarded as a source of external pressure Pext
 
P(water) = 1000 * 9.8 * 1.2 = 12 kPa.
 
Pressure p(int) exerted by the Mercury column itself is found in the same manner.
P int = 13.6 * 1000 *9.8* 0.3
= 40 kPa
 
Hence total pressure at the bottom is 52 kPa
 
Regards
Arun (askIITians forum expert)